Main INDEX
 Monthly INDEX
 PREV
 NEXT
User name x. jiang
Log entry time 23:53:50 on March 3,2000
Entry number 35336
keyword=shift activity, swing March 3rd.
16:00  start shift, waiting for one pass beam.
17:50,  beam ready.
           put in c12 target for check out.
17:53  c12 test run, beam position is changing 
17:57  beam off. due to injector work.
18:22  beam back. Joe Mitchell  preparing for e-p.
 18:30 did harp scan run-560, with raster off. check with spot.
           MCC check beam lose, indicating raster problem, take c12 target                out for check .
18:55  c12 in beam again
 19:20 set raster size. 4.0 by 4.0 mm by spot ( asked  MCC 1.5 by 5.0) 
           take harp scan run562
 19:30 start run 2528, he-3 elastic in e-arm, 5uA beam.
           prescal factor 1,1,1,10,1
           raw scaler rate 1.7K, 88, 180,134,0.
           dead time 21%
           check OTR reading , saved in halog.
           Found that OTR was not with proper mask.
20:00 e-p measurement
 21:00-22:00 harp many scan, to make sure beam spot size
          is less than 200um in both x and y.
22:00  Mcc cycle the qurds in hall-A line after beam tuning.
22:22 c12 elastic run, raster off,
          run 2531, high dead time, reset coda.
 22:30-23:25, c12 run, raster off and raster on. 
 23:25 arun start arc energy measurement.